OpenStudy (anonymous):

((3k^2+1k-2)/(4k-2)) divided by (k^2+3k+2)/(2k^2+5k-3) tells me show all work.... and Use complete sentences to explain your process.

OpenStudy (anonymous):

Hint: First factor all of your polynomials, and then recognize that \[\Huge \frac{\left(\frac{a}{b}\right)}{\left(\frac{c}{d}\right)}=\frac{ad}{bc}\]This should aid in your simplification.

OpenStudy (anonymous):

?

OpenStudy (anonymous):

i have no clue what to do

OpenStudy (anonymous):

Taking analogy to what you stated: \[a=3k^2+k-2\]\[b=4k-2\]\[c=k^2+3k+2\]\[d=2k^2+5k-3\]

OpenStudy (anonymous):

that is all right

OpenStudy (anonymous):

lets factor those... \[\large a=(k+1)(3k-2)\] \[\large b=2(2k-1)\] \[\large c=(k+1)(k+2)\] \[\large d=(k+3)(2k-1)\]

OpenStudy (anonymous):

(3k 2 +k−2/4k−2) divided by (k 2 +3k+2)/2k 2 +5k−3

OpenStudy (anonymous):

3k^2+k-2 is A

OpenStudy (anonymous):

b= 4k-2

OpenStudy (anonymous):